OpenStudy (anonymous):

MULTIPLY: k^2+2k-15 k^2+4k+3 ---------- X ---------- k^2+k-6 k^2+6k+5

OpenStudy (anonymous):

k-3/k+2

OpenStudy (anonymous):

(k+5)(k-3) (k+3)(k+1) __________ x ______________ (k+3)(k-2) (k+3(k+2) =(k+5)(k-3)(k+3)(k+1) _____________________ (k+3)(k-2)(k+3)(k+2) =(k+5)(k-3)(k+1) ________________ (k+3)(k-2)(k+2)